OpenStudy (anonymous):

the measure of the smallest angle of a triangle is 1/4 the measure of the largest angle. the measure of the second angle is 72 degrees less then the measure of the largest of the angles of the triangle. What is the largest angle, second angle, smallest angle.

OpenStudy (anonymous):

when trying to set this up i got .25x+(x-72)=180 at first is this good so far?

OpenStudy (anonymous):

don't forget about the large angle

OpenStudy (anonymous):

you have to add x as well

OpenStudy (anonymous):

for largest angle i got 94.5

OpenStudy (anonymous):

check your work

OpenStudy (anonymous):

what do you mean add x?

OpenStudy (anonymous):



OpenStudy (anonymous):

=180

OpenStudy (anonymous):

so to try to solve it would i set up equation as x/4(x-72)+x=180?

OpenStudy (anonymous):

you add the angles above = 180

OpenStudy (anonymous):

you got the right equation if you accidentally forgot the +

OpenStudy (anonymous):

ok thanks let me try plug everything back in then

OpenStudy (anonymous):

so .25(x-72)+x=180 =.25x-18+x=180 =1.25x-18=180?